What rate of change of current in a solenoid having self-inductance 19.4 mH produces a self-induced emf of 100 mV in it?
1
Expert's answer
2020-03-06T10:24:37-0500

"\lvert\Epsilon_s\rvert =L\frac{dI}{dt}"


"\frac{dI}{dt}=\frac{\lvert E_s\rvert }{L}=\frac{0.1}{0.0194}=5.15 A/s"